Fermi National Laboratory


Accelerator Update

Friday May 7
The day shift began with store #3471 in the Tevatron and colliding. The Pbar stack was at 76 mA. SY120 was turning on. Recycler was doing studies and MiniBooNe was taking beam.

At 9:57 AM MTest is running in 120 GeV mode.

At ~10:31 I:605A & I:605B faulted off and caused the MI-RF sector vacuum valves to close. The valves were opened and beam was back in 5 minutes.

At 14:05, we lost the Tevatron store when T:QDE2 tripped.

CDF and D0 were given 1 hour controlled accesses while we investigate and recover. This would allow more stacking time too.

By 15:30, D0 was out of access, but the work on T:QDE2 continued.

CDF finished their access by 16:11.

The work on T:QDE2 was finished by 16:30. They had to replace a circuit board.

By 17:05 the Tevatron was in pre-shot setup performing a dry squeeze.

Shot setup began at 17:30.

While loading final protons at 18:45, the Tevatron quenched at F32U. This was caused by loading protons on top of other protons already in the Tev.

This problem was resolved and the Tevatron recovered. The Tev was ramping again by 19:30 and was put into shot setup at 20:42.

Store 3475 was colliding with an initial luminosity of 46.8e30 by 22:30. The initial luminosity was a little lower because a MI RF vacuum burst caused us to lose about 18 mA.

Stacking resumed shortly thereafter and MCenter was secured for beam.

Saturday May 8
The midnight shift began with store 3475 in the TeV. PBar was stacking. MiniBooNe was taking beam. MCenter was starting up.

At 02:26 MCenter made a controlled access and dropped their interlocks.

At 04:37, operations went out and re-secured the area. Beam to MCenter was established when they got back.

At ~ 09:48 a problem developed with S:HT201. A power supply expert was called in to investigate and repair it.

By 11:53 it was repaired and beam was back to switchyard and the beamlines.

A problem with autotune was discovered and MiniBooNe had to be tuned manually until it was resolved around 15:00.

At ~13:00 more events were added to the timeline to assist in a Recycler/Main Injector study. This caused some timing problems, and it took a couple hours to resolve them.

At 15:20 a few Linac quad power supplies started acting up and held off beam. Operators investigated and replaced control cards. Beam returned at 15:50.

Booster RF station 16 tripped a couple of times over the next hour.

By 19:45 the RR/MI studies were complete. S:QP3 in switchyard tripped off just before this and a power supply tech had to be called for repairs.

At 22:44 the store was coming to an end. The experiments ran down their equipment and the store was terminated at 23:10.

By 23:45 the Tevatron was in shot setup again.

Sunday May 9
The midnight shift began with Operations working on the shot to the Tevatron.

Operations had the store #3477 in the TeV and colliding with an initial luminosity of 57.1e30 by 1:20 AM. Pbar returned to stacking

Beam was re-established to MCenter by 01:40.

At 05:40 Tevatron RF station #8 tripped off due to a water heater. The store stayed in the Tev.

At ~06:00 abort gap losses were up and the collimators were adjusted to reduce them.

At 08:05 a Booster vacuum burst caused two stations to trip off. They reset and came back on.

At 10:04 the abort gap losses were creeping up again and the collimators had to be moved again to reduce them.

Booster RF stations 15 and 16 tripped, but were turned back on.

Linac RF station 2 tripped, but it also reset and was turned back on.

At 21:32 Booster RF station 18 tripped. It was reset and turned back on.

MCenter went in on controlled access for about hour.

At 23:53 the MCR started to warm up. The duty mechanic found a pump problem and switched to the backup pump. The MCR started to cool down.

Monday May 10
The midnight shift began with the TeV in a store, Pbar stacking, and with MiniBooNe taking beam. MCenter was also taking beam.

At 02:15 Booster RF station 16 tripped again, was reset and turned back on.

At ~03:25 the MiniBooNe permits dropped. It was due to a power supply which was turned back on.

At ~04:00 stacking was stopped and SY120 disabled for shot setup. The store was terminated shortly thereafter.

At 04:44 shot setup was in progress.

At ~05:08 things started to fall apart. The Tev BLM was pulling the abort without beam. This was tracked to a temperature problem due to a bad AC unit. The Main Injector lost its power supply permits. An AC contactor appeared to be the problem. A power supply expert had to be called in to investigate.

The day shift began with experts working on the Main Injector power supply problem. No stacking can be done and shot setup is on hold until this can be repaired.

Today's Plan
Repair the Main Injector power supply problem and go into shot setup. Depending on the outcome of this store, Tevatron studies may be performed.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/10/2004   email Fermilab